A mixture initially contains A, B, and C in the following concentrations: [A] = 0.450 M , [B] = 1.00 M , and [C] = 0.700 M . The following reaction occurs and equilibrium is established:

A+2B⇌C

At equilibrium, [A] = 0.260 M and [C] = 0.890 M . Calculate the value of the equilibrium constant, Kc.
Answer #1

ICE Table:

Given at equilibrium,

[A] = 0.26

0.45-1x = 0.26

x = 0.19

Equilibrium constant expression is

Kc = [C]/[A][B]^2

Kc = (0.7+1x)/(0.45-1x)(1.0-2x)^2

Kc = (0.7+1*0.19)/(0.45-1*0.19)(1.0-2*0.19)^2

Kc = 8.90

Answer: 8.90